find the center and radius of the circle given by x(squared)- 4x + y(squared) -6y + 12 = 0
Answer by Nate(3500)

x^2 - 4x + y^2 - 6y = -12
(x - 2)^2 + (y - 3)^2 = -12 + 4 + 9 = 1
(x - 2)^2 + (y - 3)^2 = 1
radius = 1
center = (2,3)